# Hydrostatic mass of galaxy clusters within some theories of gravity

### v3

@@ Abstract
The mass of galaxy clusters (GCs) can be determined by calculating the hydrostatic equilibrium equation. In this work, we derive the hydrostatic mass of GCs within Eddington-inspired Born-Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and modified emergent Newtonian gravity (MENG) with generalized uncertainty principle (GUP) correction. We apply the formulations on the masses of 10 GCs. We compare our results with the Newtonian mass of GCs. Within a regime, we get an insight that all formulations could match the Newtonian mass. Thus, the impact of the modified theories of gravity used in this work can be neglected in this regime. The noteworthy impact starts if we set $\kappa=5\times 10^{40}$ for EiBI theory, $\Upsilon=-0.1655\times 10^{69}$ for BHG, and $\beta_{0}=-1.656\times 10^{110}$ for MENG. We also compare our results from EiBI theory and BHG with the baryonic masses $M_{bar}$ of the GCs. A better linear fit is achieved by EiBI theory with $\kappa_{0}=5.80\times 10^{40}$, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. This leads us to the fact that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.
@@ 2.1 Hydrostatic mass of GCs within EiBI theory
$M_{EiBI}(r)=-\left[3,7\times 10^{-4}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)+\frac{r\kappa\rho}{4M_{\odot}}\left(\frac{d\ln\rho}{d\ln r}\right)\right]\>\mathscr{M}_{\odot}.$ (6)
@@ 2.2 Hydrostatic mass of GCs within BHG
$M_{BHG}=\left[{-3,7\times 10^{-4}}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right)\right]\mathscr{M}_{\odot}.$ (13)
@@ 3 Numerical results
Figure 1: The linear fit of the relation between hydrostatic masses and the obseved baryonic mass $M_{bar}$ of 10 GCs. (a) $M_{EiBI}$ with $\kappa=5.80\times 10^{40}$. (b) (c) $M_{BHG}$ with $\Upsilon=1.65\times 10^{71}$.
@@
(kpc) ($\mathscr{M}_{\odot}$) $\beta_{0}=-1.656\times 10^{110}$ $\kappa=5\times 10^{40}$ $\Upsilon=-0.1655\times 10^{69}$
@@
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [19], [43], and [36]. In Ref. [19], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [43], $\kappa=0,2$ could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [19], polytropic star in Ref. [43], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass. However, although $\beta_{0}$ is a free parameter, setting its value to produce a significant impact observable on astronomical scales is not appropriate, as $\beta_{0}$ originates from quantum corrections.
@@
On the other hand, we also compare our results with the baryonic mass of GCs shown in Ref. [41]. Now consider
@@
The linear fit of the EiBI theory with $\kappa=5.80\times 10^{40}$ gives the slope of $0.126\pm 0.086$, while the BHG with $\Upsilon=1.65\times 10^{71}$ gives the slope of $0.023\pm 0.119$. The linear fit of EiBI theory is shown in Fig. 1(a), while the linear fit of BHG is shown in Fig. 1(b). The slope of the results from EiBI theory is closer to unity compared to that of BHG, indicating that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs, under the same assumption used in Ref. [41] of no dark matter. However, these slope values remain significantly different from unity, suggesting that neither EiBI theory nor BHG fully resolves the mass discrepancy. The exploration of alternative modified gravity theories with appropriate parameters to address the mass discrepancy problem is deferred to future work.
@@ 4 Concluding remarks
In this work, we have revisited the derivation of MENG with GUP correction derived in Ref. [19], and derive the hydrostatic mass of GCs within MENG with GUP correction. We also derive the hydrostatic mass of GCs within EiBI theory and BHG. We applied all formulations that we have revisited and derived on 10 GCs and compared the results with the Newtonian mass obtained from Ref. [3]. The impact of the theories of gravity used in this work starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG. If the values of the free parameters are set closer to zero than the values mentioned before, the impact is negligible, which means that one can safely use the Newtonian formulation to calculate the hydrostatic mass of GCs.
We compare our results of EiBI theory and BHG with $M_{bar}$. A better linear fit is achieved by EiBI theory with $\kappa_{0}=5.80\times 10^{40}$, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. This leads us to the fact that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.
@@ Acknowledgement
We are sincerely grateful to Prof. Anto Sulaksono and Idrus Husin Belfaqih for precious discussions and comments. MLP acknowledges Indonesia Endowment Fund for Education Agency (LPDP) for financial support.

### v4

@@ Abstract
The mass of galaxy clusters (GCs) can be determined by calculating the hydrostatic equilibrium equation. In this work, we derive the hydrostatic mass of GCs within Eddington-inspired Born-Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and modified emergent Newtonian gravity (MENG) with generalized uncertainty principle (GUP) correction. We apply the formulations on the masses of 10 GCs. We compare our results with the Newtonian mass of GCs. Within a regime, we get an insight that all formulations could match the Newtonian mass. Thus, the impact of the modified theories of gravity used in this work can be neglected in this regime. The noteworthy impact starts if we set $\kappa=5\times 10^{40}$ m2 for EiBI theory, $\Upsilon=-0.1655\times 10^{69}$ for BHG, and $\beta_{0}=-1.656\times 10^{110}$ for MENG. We also compare our results from EiBI theory and BHG with the baryonic masses $M_{bar}$ of the GCs. A better linear fit is achieved by EiBI theory with $\kappa=5.80\times 10^{40}$ m2, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. This leads us to the fact that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.
@@ 2.1 Hydrostatic mass of GCs within EiBI theory
$M_{EiBI}(r)=-\left[3,7\times 10^{-4}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)+\frac{r\kappa\rho}{4M_{\odot}}\left(\frac{d\ln\rho}{d\ln r}\right)\right]\>\mathscr{M}_{\odot}.$ (6)
@@ 2.2 Hydrostatic mass of GCs within BHG
$M_{BHG}=\left[{-3,7\times 10^{-4}}\>T\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right)\right]\mathscr{M}_{\odot}.$ (13)
@@ 2.3 Hydrostatic mass of GCs within MENG
Figure 1: The linear fit of the relation between Newtonian hydrostatic masses and the observed baryonic mass $M_{bar}$ of 10 GCs.
@@ 3 Numerical results
Figure 2: The linear fit of the relation between hydrostatic masses and the observed baryonic mass $M_{bar}$ of 10 GCs. (a) $M_{EiBI}$ with $\kappa=5.80\times 10^{40}$ m2. (b) $M_{BHG}$ with $\Upsilon=1.65\times 10^{71}$.
@@
(kpc) ($\mathscr{M}_{\odot}$) $\beta_{0}=-1.656\times 10^{110}$ $\kappa=5\times 10^{40}$ m2 $\Upsilon=-0.1655\times 10^{69}$
@@
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$ m2 for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [19], [43], and [36]. In Ref. [19], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [43], $\kappa=0.2$ m2 could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we realize that the objects used are white dwarf in Ref. [19], polytropic star in Ref. [43], and brown dwarf in Ref. [36]. Compared to GCs, those objects have different structures and EOS. So it is clear that GCs have their own values of the free parameters (i.e. $\beta_{0}$, $\kappa$, and $\Upsilon$) to conduct some impact on the mass. However, although $\beta_{0}$ is a free parameter, setting its value to produce a significant impact observable on astronomical scales is not appropriate, as $\beta_{0}$ originates from quantum corrections.
@@
On the other hand, we also compare the Newtonian mass of GCs and our results with the baryonic mass of GCs shown in Ref. [41]. The Newtonian mass, derived from the hydrostatic equilibrium equation, corresponds to the Tolman-Oppenheimer-Volkoff (TOV) equation in GR. The GC mass obtained from the TOV equation has been studied in Ref. [3], where numerical results indicate that the GR correction is negligible. Therefore, the Newtonian GC mass can be considered equivalent to the GR mass in the Newtonian limit. Furthermore, if the parameter sets of the EiBI theory, BHG, and MENG are set to zero, the formalisms consequently recover to the Newtonian mass. Now consider
@@
The linear fit of the Newtonian mass shown in Fig. 1 gives the slope of $0.112\pm 0.012$. The linear fit of the EiBI theory with $\kappa=5.80\times 10^{40}$ m2 gives the slope of $0.126\pm 0.086$, while the BHG with $\Upsilon=1.65\times 10^{71}$ gives the slope of $0.023\pm 0.119$. The linear fit of EiBI theory is shown in Fig. 2(a), while the linear fit of BHG is shown in Fig. 2(b). The slope of the results from EiBI theory is closer to unity compared to that of BHG, indicating that EiBI theory is more effective than BHG in alleviating the mass discrepancy between hydrostatic mass and baryonic mass in GCs, under the same assumption used in Ref. [41] of no dark matter. Moreover, the slope of the results from the EiBI theory is closer to unity compared to those from the Newtonian mass. Nevertheless, the standard error of the linear fit in the EiBI theory is larger than that of the Newtonian mass. However, these slope values remain significantly different from unity, suggesting that neither EiBI theory nor BHG fully resolves the mass discrepancy. The exploration of alternative modified gravity theories with appropriate parameters to address the mass discrepancy problem is deferred to future work.
@@ 4 Concluding remarks
In this work, we have revisited the derivation of MENG with GUP correction derived in Ref. [19], and derive the hydrostatic mass of GCs within MENG with GUP correction. We also derive the hydrostatic mass of GCs within EiBI theory and BHG. We applied all formulations that we have revisited and derived on 10 GCs and compared the results with the Newtonian mass obtained from Ref. [3]. The impact of the theories of gravity used in this work starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ m2 for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG. If the values of the free parameters are set closer to zero than the values mentioned before, the impact is negligible, which means that one can safely use the Newtonian formulation to calculate the hydrostatic mass of GCs.
We compare our results of EiBI theory and BHG with $M_{bar}$. A better linear fit is achieved by EiBI theory with $\kappa_{0}=5.80\times 10^{40}$ m2, which gives the slope $\mathcal{M}$ of $0.126\pm 0.086$. This value is closer to unity than the one of BHG. Moreover, the slope of the results from the EiBI theory is closer to unity compared to those from the Newtonian mass. However, the standard error of the linear fit in the EiBI theory is larger than that of the Newtonian mass. This indicates that, compared to BHG, EiBI theory is more effective in reducing the mass discrepancy between the hydrostatic mass and baryonic mass in GCs. Nevertheless, neither EiBI theory nor BHG completely addresses the mass discrepancy problem.
@@ Acknowledgement
We are sincerely grateful to Prof. Anto Sulaksono and Idrus Husin Belfaqih for precious discussions and comments. We sincerely thank the reviewers for their thorough evaluation of our manuscript. Their insightful comments and constructive suggestions have significantly improved the clarity, accuracy, and overall quality of our work. We deeply appreciate their time and effort in providing such valuable feedback. MLP acknowledges Indonesia Endowment Fund for Education Agency (LPDP) for financial support.